# Importance of accounting for student identities and intersectionality for creating equitable and inclusive physics learning environments

### v1

@@ III.1 Pathway to Majoring in Physics
In this section, we discuss the experiences that influenced these women’s decision to major in physics in college, e.g., family’s role, and positive high school experiences, while also discussing their negative high school experiences.
III.1.1 Family’s role in providing support
@@
All of the women described several ways in which their family provided support to them, e.g., providing general support, academic support, and exposure to science at an early age. These positive experiences with family as described below were very important in putting these women on the path to majoring in physics and potentially creating a science identity.
@@
Family providing academic support: Many women described how having family members who are physicists, who majored in physics, or had STEM degrees, which required taking physics courses, have been able to provide academic support in their physics journey through having first hand experiences with related disciplines themselves.
@@
Cathy describes how her sister, who is doing her Ph.D. in physics, helped her learn physics and provided encouragement to major in physics: “I kind of like learned a lot about it from her and she encouraged me to go into it…I think that she made it a lot easier for me to go into it…she had a lot of good advice and…just kind of made the whole thing like pretty accessible.” Having someone directly in the field was very beneficial to Cathy, making physics ‘easier’ to get into and ‘pretty accessible’ to consider as a college major. Receiving advice from an insider, someone who already went through the process, may have been especially beneficial as that advice may have been specific to physics because even within different STEM disciplines, the established norms and cultures can be different..
Similarly, Francis explains how one of her older sisters who majored in engineering has provided academic support: “whenever I had issues with like homework and wanted help, I would always…go to her and she would help me out which I guess probably helped a lot just knowing that I had someone who could help…” Again, having family who can help with homework is very beneficial, especially when her older sister already took physics classes and hence was familiar with the material.
@@
Chloe also describes how her parents being physicists have influenced her decision to major in physics: “I’ve been interested in physics, most of my life, probably because both of my parents studied physics and are physicists and all their family friends are also physicists. I’ve always been surrounded by it, so I think I almost had an advantage in that way, in which seeing actual physicists and particularly female physicists-it had a big influence on me and I’ve always kind of wanted to pursue physics, particularly astronomy…” Being surrounded by physicists was beneficial to her, e.g., not only is her self-efficacy likely to get a boost particularly because her mother is a physicist but close family ties with physics possibly introduced her to various topics in physics and astronomy early on and provided insights as to what a career in physics may look like. In particular, seeing female physicists may have been especially beneficial in that Chloe was seeing people with a shared gender identity pursuing and being successful in a field that she had interest in.
@@
Ruby has an uncle who is a physicist. She explains that during a previous semester, her uncle provided support by recommending books: “last semester, when I was taking modern physics I would talk to him about it and he like recommended this book on matrices and commuting principles and I-it was a little out of my understanding, but it was still really interesting.” Although the books that Ruby’s uncle gave her were advanced, having someone like her uncle who supports her learning and interests can foster interests that could firm up the decision to majoring in physics.
@@
Esme describes how her dad instilled the importance of education into her: “He works in construction and…since he was born in Mexico, he couldn’t get an education here in the US, so he kind of always projected education is important. It was the number one thing I had to focus on and he was always into engineering and all that stuff and because we were so close, that’s how I got into physics. I’ve always been into astronomy. So he had always…encouraged me to go into whatever I liked and I’ve always been very STEM oriented because of him.” Thus, Esme’s dad encouraged her academic endeavours and his interest in engineering influenced her own interest in physics.
@@
We see the role family has had in supporting these women majoring in physics at this liberal arts college. Not only did family provide academic support, if they had careers in STEM, but also provided opportunities for these women to interact and learn science in contexts that were applicable to their jobs or pursued scientific activities for fun, e.g., Ruby’s puzzles with her grandfather.
III.1.2 Positive high school experiences
Many if not all of these women took a physics class during high school. We recount their positive experiences in their high school physics class. Positive experiences include perceptions of and/or interactions with their teachers.
@@ III.1.2 Positive high school experiences
Esme attended a public school and shares that she was part of a NASA program during her junior year. “I got to go to NASA for a week, all expense paid trip and I would be there and…we did like a mock Mars mission and I got to work with actual engineers and they would give us lectures and give us tours around NASA, it was very great opportunity…” This positive experience may have solidified her interest in physics and astronomy. It allowed her to participate in astronomy by working with NASA engineers and learning about astronomy.
@@
Francis attended an all girls high school in the UK. She summarizes, “because it was just all girls…I didn’t feel out of place and I guess that helped, yeah I guess I’ve always had quite nice physics teachers.” She explains that when it came to college applications, her physics teachers were, “always very willing to help with applications and interviews and that kind of thing, because in the UK, when you apply, you apply [for] one degree…some places, will have interviews and they’ll ask you questions about that, so I remember, I did a [mock] interview with one of my physics teachers…and they would ask physics questions and solve it and talk through how I solved it and they would definitely [be] very helpful with that.” Thus, Francis’ physics teachers played a very important role in helping her prepare for college, in addition to encouraging her and other students to apply for college.
@@
Paulina recalls her interest in physics and taking several classes during high school. Through this class she was able to have a lot of hands-on experience in labs and was able to have one-on-one conversations with her teacher: “I became really interested in the class and I would stay after [to] chat with him, ask questions, help other students and I think at that point, I just kind of viewed it as a hobby of mine. I didn’t really expand on it too much and so after that I proceeded to take AP physics C… electromagnetism and mechanics and that was also very interesting to me, I think it was one of the first experiences that I had where I felt truly challenged in a class, but also in a very positive way, I felt I was getting something out of the education I was receiving and it was very interesting, especially because I could visually see the ways in which physics and I interacted a lot in my everyday life. And that was what really first got me into physics.” It is possible that having these one-on-one conversations with her high school teacher was beneficial and helped her see physics as an interesting field to pursue, and potentially a field in which she can have a career instead of pursuing it only as a hobby. These conversations may have also encouraged her to continue challenging herself, such as by taking the AP Physics exam. Her perception of physics throughout the interview was very positive and she found connections between physics she learned in class and physics she interacted with in everyday life.
@@
Chloe attended a private school which she says was known for being a “competitive college prep school where there are many students who had been like training in math Olympiads and science competitions their whole lives.” She had overall mixed experiences during high school, but in this section we will comment on her positive experiences. Despite the negative competitive culture, she says, “…it was inspiring to be around people who were so driven academically, so I think I grew academically always [like] I wouldn’t have if I hadn’t been surrounded by these people who are just really focused on getting into college.” The fact that students at her high school were driven and focused on college education encouraged her to also focus on her academic trajectory.
@@
Karina also attended a private high school and had mixed experiences. Her positive experiences were mainly in regard to her high school teachers. When asked about potential gendered experiences in high school, she says, “Both teachers were male, but they were very supportive, and I don’t think I ever felt treated unfairly because of my gender…I always felt they treated me like I was smart and capable and deserved to be there…I can definitely remember for that advanced physics class my senior year, there was a big gender disbalance, and I think even if people are treating you fairly, if you feel like one in the minority, it does affect your experience.’ Despite there being more men than women in her high school physics class in her senior year, Karina still felt supported by her teachers. She also adds an important fact that having an underrepresented identity can affect a student’s experience even if others are treating them “fairly”.
@@
III.1.3 Negative experiences in high school
@@
When asked about her high school experiences, Chloe says: “I think the competitive culture was quite toxic I’d say.” In regard to when students applied to college, she recalls, “I remember when the early decisions came out, it became universally banned to talk about where you had gotten in because there is so much competition amongst students and so many people would feel bad [but] everybody knew where everybody had gotten in because it was the only topic of gossip our senior year.” It seems like students in Chloe’s senior class were fixated on their college status and students did not communicate with each other positively due to tensions related to college admissions. Chloe, who described herself as an introvert, also recalls negative experiences in her math class, in which, “sometimes people would talk and when I would talk nobody would hear me, but when some of the quiet boys would talk, everybody would listen to them.” She reflects upon and explains that her being ignored might be a result of students stereotyping her, e.g., male students assuming that she does not have anything substantial to contribute to conversations.
@@
Karina feels guilty about going to a private high school, she says: “If I think morally about private school education, I don’t believe in it, so I felt guilty about having that opportunity. And it was a pretty wealthy school.” Having access to better educational resources was a point of guilt and contention that she points out several times in her interview. She also shares that the culture of her high school was discouraging: “I got kind of discouraged by it. It was kind of like a college prep type environment, where the sort of expectation was to go to an Ivy league school or that’s a failure and to have like-I felt like everyone around me was thinking a lot about GPA and not really caring about the learning so I got a little discouraged…”. Thus, in Karina’s high school, the culture fostered high expectations for students to earn good grades to attend top universities as opposed to encouraging them to focus on learning, which made her feel discouraged.
@@
Esme talks about her physics teacher and how she did not like the priorities he had for students: “I really didn’t like my physics class… [My physics teacher] really didn’t do a good job…We would just work on one problem, and he would just kind of vaguely explain the physics of it, but we mostly worked on trying to pass the AP [exam] rather than understand the concepts of physics. So, I didn’t really like that class. It almost made me really quit physics, because…I didn’t want to do just problems and not understand.” Esme did not feel like she was learning physics, but rather working to pass the AP exam. Since the focus of the physics class was not to understand concepts, she did not like it and if it was not for a very positive NASA experience, she may not have chosen a physics and astronomy major in college. She also mentions that her high school teacher called out only a few students who he thought would be the only ones to pass the AP exam if they try hard enough (implying that others were not capable).
@@
When reflecting on how being a woman affected her interactions with her high school teacher, she says: “but in terms of being a woman, my teacher was very strange, he… had a weird relationship with the girls in my class.” She clarifies that “he would always compliment them and…just be kind of unprofessional, like not professional, at one point I remember he made a Tinder [account] and he wanted help so some of the girls in my class were helping him.” For reference, Tinder is an online dating app. Esme also noted that his inappropriate behavior with female students made the environment weird and uncomfortable.
@@
Her interview also points to the fact that Esme had negative interactions with her male peers: “They knew they also wanted to go into engineering and physics and they were always about the money and always like, Yeah, I got this, like they were very arrogant and yet they would kind of…mooch off of my work. They always tried to take my work…If we had partners, I usually did my portion [of] work, and I would try to make them do it, but sometimes they wouldn’t and it would affect my grade so I would have to do everything.” Thus, we see that Esme’s high school male peers took advantage of her work and sometimes channelled it as their own or they would not contribute and leave her to do things on her own. This account corresponds to findings in another study where male students dominated group work [46].
@@
When asked why Esme found it difficult to communicate with her male peers, she says: “It really seemed like they didn’t respect me. I always thought maybe it was just my personality…Personally, I’m very flexible…I usually put up with things, but I have a limit and they would reach that limit many times.” Despite the fact that she describes herself as having a reasonable amount of tolerance for her male peers’ behavior, she thinks that her personality might be the reason why the interactions with them often led to friction. This is important to note as it may be a reflection of the masculine STEM culture and how it can make women think that they are the problem when it comes to negative gender dynamics and interactions with male peers. She also noted that she made it a point to avoid her male peers. When asked about the gender distribution of her peers in her high school classes, she says that “there were a little more men than women but I honestly can’t say how many more because… I really would try to not interact with them unless they were someone I knew.” This avoidance might have been stressful but it may have been a survival mechanism in that the avoidance may prevent her from experiencing any biases or microaggressions from her peers first hand.
@@
Ruby describes her high school physics experiences as: “I liked high school physics, a lot. I did well in it, but I never felt like it was something that I could go into honestly before college. [At] my high school, no one ever talks about becoming a physics major and no one ever even tells you that that’s necessarily an option, and until I got to college, I really didn’t think that that was something I was even capable of if that makes sense.” Although she liked her physics class and her high school teachers, she never felt encouraged to pursue physics as a college degree. Not having teachers encourage students to pursue physics could be extremely discouraging especially for students from traditionally marginalized groups. The lack of encouragement could also be interpreted as teachers not thinking students are capable of succeeding in a physics major, which may have contributed to Ruby feeling this way.
@@
When recalling her high school physics teacher, Ruby says: “I definitely noticed that in high school, my teacher would always-even if my hand was up or if another girl’s hand is-that he would always call on the male students before the female students.” Recollecting that her high school teacher showed gender bias, she further says, “I do think there was still this bias on his side of thinking, Oh, the guys are able to do this more than the girls. Well, oh, the girls will need your [more] help with laboratory experiments than their male counterparts. And there definitely was that tone under the setting. But it wasn’t obviously apparent.” Ruby seemed to catch on to these biases which implicitly reflected that women may not be as good as men in physics or at physics lab. Despite the possibility that these biases may have had lasting experience on her, Ruby says, “I sort of just brushed it off because I didn’t want that [high school teacher’s gender bias] to get in the way of me doing something that I liked.” It is great that she didn’t let a teacher’s biased comments affect her self-efficacy and prevent her from pursuing physics but it could have affected other women in her class.
@@
III.2 Interactions with college peers
In this section, we describe the experiences these women had with their peers in college physics courses. These experiences include positive and negative interactions and perceptions.
III.2.1 Positive interactions and perceptions

### v2

@@ IV.1 Pathway to Majoring in Physics
In this section, we discuss the experiences that influenced the women’s decision to major in physics in college, e.g., family’s role, and positive high school experiences, while also discussing their negative high school experiences.
IV.1.1 Family’s role in providing support
@@
All of the women described several ways in which their family provided support to them, e.g., providing general support, academic support, and exposure to science at an early age. These positive experiences with family as described below were very important in putting the women on the path to majoring in physics and potentially developing a science identity.
@@
Family providing academic support: Many women described how family members supported them in their physics journey through having first hand experiences with related disciplines themselves. For instance, some described that family members were physicists, majored in physics, or had STEM degrees, which required taking physics courses.
@@
Cathy describes how her sister, who is doing her Ph.D. in physics, helped her learn physics and provided encouragement to major in physics: “I kind of like learned a lot about it from her and she encouraged me to go into it…I think that she made it a lot easier for me to go into it…she had a lot of good advice and…just kind of made the whole thing like pretty accessible.” Similarly, Francis explains how one of her older sisters who majored in engineering has provided academic support. For example, she is able to go to her sister for homework help. Here we see that having someone directly in the field was very beneficial to Cathy and Francis. In addition, receiving advice from an insider, someone who already went through the process, may have been especially beneficial as that advice may have been specific to physics because even within different STEM disciplines, the established norms and cultures can be different.
@@
Chloe also describes how her parents being physicists have influenced her decision to major in physics. She says, “I’ve always been surrounded by [physics], so I think I almost had an advantage in that way, in which seeing actual physicists and particularly female physicists-it had a big influence on me and I’ve always kind of wanted to pursue physics, particularly astronomy…” Being surrounded by physicists was beneficial to her, e.g., not only is her self-efficacy likely to get a boost particularly because her mother is a physicist but close family ties with physics possibly introduced her to various topics in physics and astronomy early on and provided insights as to what a career in physics may look like. In particular, seeing female physicists may have been especially beneficial in that Chloe was seeing people with a shared gender identity pursuing and being successful in a field that she had interest in.
@@
Esme describes how her dad instilled the importance of education into her: “He works in construction and…since he was born in Mexico, he couldn’t get an education here in the US, so he kind of always projected [that] education is important. It was the number one thing I had to focus on and he was always into engineering and all that stuff and because we were so close, that’s how I got into physics. I’ve always been into astronomy. So he had always…encouraged me to go into whatever I liked and I’ve always been very STEM oriented because of him.” Thus, Esme’s dad encouraged her academic endeavours and his interest in engineering influenced her own interest in physics.
@@
We see the role family has had in supporting the women majoring in physics at this liberal arts college. Not only did family provide academic support, if they had careers in STEM, but also provided opportunities for the women to interact and learn science in contexts that were applicable to their jobs or pursued scientific activities for fun, e.g., Ruby’s puzzles with her grandfather.
IV.1.2 Positive high school experiences
Many of the women took a physics class during high school. We recount their positive experiences in their high school physics class. These positive experiences capture a range of interactions and perceptions of their peers and/or teachers. For example, they noted that the teachers were supportive and peers were driven, which inspired them.
@@ IV.1.2 Positive high school experiences
Esme attended a public school and shares that she was part of a NASA program during her junior year. “I got to go to NASA for a week, all expense paid trip and I would be there and…we did like a mock Mars mission and I got to work with actual engineers and they would give us lectures and give us tours around NASA, it was very great opportunity…” This positive experience may have solidified her interest in physics and astronomy by allowing her to work with NASA engineers and learning about astronomy.
@@
Francis attended an all girls high school in a European country. She didn’t feel out of place because, “it was just all girls…” She also describes having “nice physics teachers.” For example, she explains that when it came to college applications, her physics teachers were, “always very willing to help with applications and interviews.” She explains that in her country of origin, when applying for a college degree, some universities will have interviews. Francis also describes how her physics teachers would help by doing mock interviews in which “they would ask physics questions and solve it and talk through how I solved it and they would definitely [be] very helpful with that.” Thus, Francis’ physics teachers played a very important role in helping her prepare for college, in addition to encouraging her and other students to apply for college.
@@
Paulina recalls her interest in physics and taking several classes during high school. Through this class she was able to have a lot of hands-on experience in labs and was able to have one-on-one conversations with her teacher: “I became really interested in the class and I would stay after [to] chat with him, ask questions, help other students and I think at that point, I just kind of viewed it as a hobby of mine. I didn’t really expand on it too much and so after that I proceeded to take AP physics C… electromagnetism and mechanics and that was also very interesting to me. I think it was one of the first experiences that I had where I felt truly challenged in a class, but also in a very positive way. I felt I was getting something out of the education I was receiving and it was very interesting, especially because I could visually see the ways in which physics and I interacted a lot in my everyday life. And that was what really first got me into physics.” It is possible that having these one-on-one conversations with her high school teacher was beneficial and helped her see physics as an interesting field to pursue, and potentially a field in which she can have a career instead of pursuing it only as a hobby. These conversations may have also encouraged her to continue challenging herself, such as by taking the AP Physics exam. Her perception of physics throughout the interview was very positive and she found connections between physics she learned in class and physics she interacted with in everyday life.
@@
Chloe attended a private school which she says was known for being a “competitive college prep school where there are many students who had been like training in math Olympiads and science competitions their whole lives.” She had overall mixed experiences during high school, but in this section we will comment on her positive experiences. Despite the negative competitive culture, she says, “…it was inspiring to be around people who were so driven academically, so I think I grew academically… I wouldn’t have if I hadn’t been surrounded by these people who are just really focused on getting into college.” The fact that students at her high school were driven and focused on college education encouraged her to also focus on her academic trajectory.
@@
Karina also attended a private high school and had mixed experiences. Her positive experiences were mainly in regard to her high school teachers. When asked about potential gendered experiences in high school, she says, “Both teachers were male, but they were very supportive, and I don’t think I ever felt treated unfairly because of my gender…I always felt they treated me like I was smart and capable and deserved to be there…I can definitely remember for that advanced physics class my senior year, there was a big gender disbalance, and I think even if people are treating you fairly, if you feel like one in the minority, it does affect your experience.” Despite there being more men than women in her high school physics class in her senior year, Karina still felt supported by her teachers. She also adds an important fact that having an underrepresented identity can affect a student’s experience even if others are treating them “fairly”.
@@
IV.1.3 Negative experiences in high school
@@
When asked about her high school experiences, Chloe says: “I think the competitive culture was quite toxic I’d say.” When students applied to college and received early decision, she recalls, “it became universally banned to talk about where you had gotten in because there is so much competition amongst students and so many people would feel bad [but] everybody knew where everybody had gotten in because it was the only topic of gossip our senior year.” It seems like students in Chloe’s senior class were fixated on their college status and students did not communicate with each other positively due to tensions related to college admissions. Chloe, who described herself as an introvert, also recalls negative experiences in her math class, in which, “sometimes people would talk and when I would talk nobody would hear me, but when some of the quiet boys would talk, everybody would listen to them.” She reflects upon and explains that her being ignored might be a result of students stereotyping her, e.g., male students assuming that she does not have anything substantial to contribute to conversations.
@@
Karina feels guilty about going to a private high school, she says: “If I think morally about private school education, I don’t believe in it, so I felt guilty about having that opportunity. And it was a pretty wealthy school.” Having access to better educational resources was a point of guilt and contention that she points out several times in her interview. She also shares that the culture of her high school was discouraging: “I got kind of discouraged by it. It was kind of like a college prep type environment, where the sort of expectation was to go to an Ivy league school… I felt like everyone around me was thinking a lot about GPA and not really caring about the learning so I got a little discouraged…”. Thus, in Karina’s high school, the culture fostered high expectations for students to earn good grades to attend top universities as opposed to encouraging them to focus on learning, which made her feel discouraged.
@@
Esme talks about her physics teacher and how she did not like the priorities he had for students: “[My physics teacher] really didn’t do a good job…We would just work on one problem, and he would just kind of vaguely explain the physics of it, but we mostly worked on trying to pass the AP [exam] rather than understand the concepts of physics. So, I didn’t really like that class. It almost made me really quit physics, because…I didn’t want to do just problems and not understand.” Esme did not feel like she was learning physics, but rather working to pass the AP exam. Since the focus of the physics class was not to understand concepts, she did not like it and if it was not for a very positive NASA experience, she may not have chosen a physics and astronomy major in college. She also mentions that her high school teacher called out only a few students who he thought would be the only ones to pass the AP exam if they try hard enough (implying that others were not capable).
@@
When reflecting on how being a woman affected her interactions with her high school teacher, Esme says: “but in terms of being a woman, my teacher was very strange, he… had a weird relationship with the girls in my class.” She explains: “he would always compliment them and…just be kind of unprofessional…” She also recounts when her teacher would ask girls in the class for dating advice. In addition to a lack of emphasis on understanding concepts, this uncomfortable environment created by her high school physics teacher may have made it difficult to focus on learning.
@@
Her interview also points to the fact that Esme had negative interactions with her male peers. For instance, she describes her peers as always thinking about “the money” because they wanted to go into engineering and physics. She says that her peers were like, “Yeah, I got this, like they were very arrogant and they would kind of…mooch off of my work. They always tried to take my work…If we had partners, I usually did my portion [of] work, and I would try to make them do it, but sometimes they wouldn’t and it would affect my grade so I would have to do everything.” Esme’s account of her male peers taking advantage of her work and leave her to do things on her own corresponds to findings from a previous study in which male students dominated group work [46].
@@
When asked why Esme found it difficult to communicate with her male peers, she says: “It really seemed like they didn’t respect me. I always thought maybe it was just my personality…I usually put up with things, but I have a limit and they would reach that limit many times.” Despite the fact that she describes herself as having a reasonable amount of tolerance for her male peers’ behavior, she thinks that her personality might be the reason why the interactions with them often led to friction. This is important to note as it may be a reflection of the masculine STEM culture and how it can make women think that they are the problem when it comes to negative gender dynamics and interactions with male peers. She also noted that she made it a point to avoid her male peers. When asked about the gender distribution of her peers in her high school classes, she says that “there were a little more men than women but I honestly can’t say how many more because… I really would try to not interact with them unless they were someone I knew.” This avoidance might have been stressful but it may have been a survival mechanism in that the avoidance may prevent her from experiencing any biases or microaggressions from her peers first hand.
@@
Ruby says she liked her physics class and although she did well in the class, she says, “I never felt like it was something that I could go into honestly before college. [At] my high school, no one ever talks about becoming a physics major and no one ever even tells you that that’s necessarily an option, and until I got to college, I really didn’t think that that was something I was even capable of if that makes sense.” Although she liked her physics class and her high school teachers, she never felt encouraged to pursue physics as a college degree. Not having teachers encourage students to pursue physics could be extremely discouraging especially for students from traditionally marginalized groups. The lack of encouragement could also be interpreted as teachers not thinking students are capable of succeeding in a physics major, which may have contributed to Ruby feeling this way.
@@
When recalling her high school physics teacher, Ruby says: “even if my hand was up or if another girl’s hand is [up]-that [my teacher] would always call on the male students before the female students.” Recollecting that her high school teacher showed gender bias, she further says, “I do think there was still this bias on his side of thinking, Oh, the guys are able to do this more than the girls. Well, oh, the girls will need your [more] help with laboratory experiments than their male counterparts. And there definitely was that tone under the setting. But it wasn’t obviously apparent.” Ruby seemed to catch on to these biases which implicitly reflected that women may not be as good as men in physics or at physics lab. Despite the possibility that these biases may have had lasting experience on her, Ruby says, “I sort of just brushed it off because I didn’t want [my high school teacher’s gender bias] to get in the way of me doing something that I liked.” It is great that she did not let a teacher’s biased comments affect her self-efficacy or prevent her from pursuing physics but it could have affected other women in her class.
@@
IV.2 Interactions with college peers
In this section, we describe the experiences the women had with their peers in college physics courses. These experiences include positive and negative interactions and perceptions.
IV.2.1 Positive interactions and perceptions